The enthalpy of solution of sodium hydroxide is -42.6 kJ.mol-1. What mass of sodium hydroxide would produce a temperature rise of 5.00 K when dissolved in 200.0 cm3 of water?


1
Expert's answer
2021-01-14T04:26:13-0500

q = msΔt

mass of solution = mass of water + mass of sodium hydroxide

"m = m_w + m_{sh}"

M(NaOH) = 40 g/mol

n(NaOH) "= \frac{m_{sh}}{40}"

1 mole of NaOH gives heat energy of 44.2 KJ

So, "q = 44200 \times \frac{m_{sh}}{40}"

The heat capacity is 4.184 J/C

"44200 \times \frac{m_{sh}}{40} = (200 + m_{sh}) \times 4.184 \times 5 \\

1105 \times m_{sh} = 4184 + 20.92 \times m_{sh} \\

1084.08 \times m_{sh} = 4184 \\

m_{sh} = 3.86 g"

Answer: 3.86 g